# Contrastive Sequential-Diffusion Learning: Non-linear and Multi-Scene Instructional Video Synthesis

### v2

@@ Abstract
Generated video scenes for action-centric sequence descriptions like recipe instructions and do-it-yourself projects include non-linear patterns, in which the next video may require to be visually consistent not on the immediate previous video but on earlier ones. Current multi-scene video synthesis approaches fail to meet these consistency requirements. To address this, we propose a contrastive sequential video diffusion method that selects the most suitable previously generated scene to guide and condition the denoising process of the next scene. The result is a multi-scene video that is grounded in the scene descriptions and coherent w.r.t the scenes that require visual consistency. Experiments with real-world action-centric data demonstrate the practicality and improved consistency of our model compared to prior work.
@@ 1 Introduction
When people perform a task with numerous intricate steps, complementing the textual step instructions with visuals enhances user experience [11, 25]. For this reason, various platforms and tools offer multi-scene videos to convey instructional textual descriptions, such as recipe instructions and do-it-yourself (DIY) projects [16].
State-of-the-art video synthesis methods demonstrate remarkable performance in generating single-scene videos [12, 26, 29, 3, 4]. Yet, only a few works address multi-scene video generation [18, 17, 31]. These methods focus on domains where a single character is central to all scenes, achieving coherence by reusing and combining visual elements across frames. However, multi-scene instructional video synthesis raises a number of challenges. First, the input is a strict sequence of actions, for which there’s a need to generate the full sequence of videos. A model should not generate just the last scene, like GILL [14], and one cannot provide the topic and let the model generate a random sequence of text-video pairs similarly to VideoDrafter [18]. Second, similar to story generation [21, 9, 23, 19], the generative model needs to determine which previous step in the sequence to ground each new scene to. Third, while existing methods are focused on video generation in which one single (typically human) character is the center of all scenes [18], instructional videos typically incorporate multiple objects instead of central characters. Hence, we argue that multi-scene instructional video synthesis requires an approach that is sequence-grounded by design, (see Figure 1).
Figure 1: CoSeD is grounded on an input sequence of step actions to synthesize non-linear, multi-scene instructional videos. (Code and data will be available after publication; videos can be accessed here)
To this end, we propose CoSeD (Contrastive Sequential Diffusion learning), a novel approach for instructional video generation. Our method generates candidate images for each step based on the textual description and latent information from previous steps. We then use a contrastive selection approach to choose the best image by evaluating it against prior step descriptions and images. Finally, we use these images to produce a video for each step, ensuring an accurate and consistent representation of the entire task sequence.
CoSeD was able to generate coherent video sequences across diverse instructional content, maintaining high fidelity and relevance in aligning language with vision. CoSeD showed a 20% improvement in human evaluations compared to existing multi-scene methods and it was preferred 68% of the time in side-by-side comparisons. Additionally, the compact size of our model enables efficient training and fine-tuning. To summarize our contributions:
@@
  • We introduce contrastive-diffusion learning over latents sampled from previous generations.
  • We enhance understanding of the role of seeds and conditioned latents in the reverse diffusion process.
@@ 2 Related Work
Various approaches have been explored to address coherence in image generation. AR-LDM [21] introduces a history-aware autoregressive latent diffusion model that incorporates information from previous steps into the diffusion model’s cross-attention mechanism to guide generation. However, achieving the reported results requires intensive training of the entire pipeline for each dataset, which includes 650 million parameters. Make-a-Story [23] incorporates the complete history of intermediate image representations (latent vectors), which may introduce noise and potentially lead to content generation based on less relevant past information.. GILL [14] fuses frozen text-only arge language models (LLMs) with pre-trained image encoder and decoder models through a mapping network, enabling multimodal capabilities like image retrieval and generation. However, this can break coherence if retrieved images don’t align with the context. SEED-LLaMA [10] integrates a visual tokenizer with a multimodal LLM to process text and images, excelling in multi-turn generation, but struggles with maintaining narrative coherence in story generation tasks.
@@
To generate long single-scene videos Blattmann [4] and Yin [30] generate sparse key frames and interpolates intermediary frames between them recursively to enhance the frame rate. Extending this idea, Stable Video Diffusion [3] creates a large dataset of annotated video clips by filtering out those with low motion or excessive text, resulting in the generation of higher-quality videos. By contrast, Lumiere [2] generates the entire video in a single pass, eliminating the need for sparse key frames and interpolation.
@@
Improvements have also been made in generating coherent multi-scene videos. Video Drafter [18] employs brute-force LLM prompting to create distinct scenes and detailed descriptions for each element. It then generates image templates, which are combined with the scene descriptions to produce the final video. Similarly, VideoDirectorGPT  [17] employs a two-stage process where GPT-4 [20] expands text prompts into detailed descriptions and ensures visual continuity by generating textual descriptions and entity layouts. Mora [31] introduces a multi-agent framework, breaking down tasks into subtasks like refining prompts [27], and generating images to create small video segments, which are then assembled [7] into a final coherent video, achieving performance comparable to closed-source models like SORA [6]. However, it relies on using only the last frame of each video segment to start the generation of the next one, which can be problematic if the current step isn’t directly related to the previous one. StoryDiffusion [32] maintains coherence across frames using a Self-Attention mechanism and a module for smooth transitions, ensuring videos faithfully depict the input text’s content. Lastly, TALC [1] enhances text-to-video (T2V) models by improving temporal alignment between video scenes and text segments, enhancing visual fidelity and narrative coherence.
@@
Building on these advancements, we develop a method that addresses the challenge of maintaining scene coherence while respecting text descriptions, all while keeping the model compact to facilitate fine-tuning across multiple domains.
@@ 3 Problem Setting
3.1 Sequential Step Dependency
@@
In our setting, we depart from the linear dependency assumption used by previous works [31] and acknowledge the possibility of a more complex and non-linear sequential structure [8]. To address this assumption, the model needs to consider not only the current step description and visual scene pair ($s_{n}$, $v_{n}$), but also the pairs from the previous steps and visual scenes, $\{(s_{1},v_{1}),\ldots,(s_{n-1},v_{n-1})\}$. This ensures coherence in the generated visual elements, maintaining consistency and accurately reflecting task progression, even when individual steps are vague or missing details.
3.2 Sequential Diffusion
Latent Diffusion Models are designed to synthesize one single image or video at a time. Our goal is to move beyond this limitation and propose a sequential-diffusion model that learns how semantic and visual dependencies should exist in a sequence of multiple-scenes. Using the Latent Diffusion Models formulation proposed by [24], the independent denoising process for each isolated step $s_{n}$ of a sequence is the direct application of the model,
@@ 3.2 Sequential Diffusion
where $z_{t}^{n}$ corresponds to the denoising iteration $t$ of the visual scene $v_{n}$, hence $z_{0}^{n}=E(v_{n})$. Formally, we wish to learn a sequence model that iteratively estimates the $v_{n}$ scene that maximizes the likelihood of the entire sequence of all previous $n-1$ steps,
@@
where $N$ is the total number of steps in a given sequence. We propose to ensure the sequential consistency through the text conditioning encoder $\tau_{\theta}(s_{i})$ and the visual denoising seed $z^{n}_{T}$. The proposed contrastive sequential-diffusion learning, Figure 2,incorporates these two methods and will be discussed in the next section.
@@ 4 CoSeD: Contrastive Sequential Diffusion
Our approach aims to find the most accurate image for generating a video that depicts a step of a task. We begin by using a decoder to better align the step description with a visual caption/prompt (Section 4.1). Then we generate candidate images based on information from previous images and the visual caption of the current step (Section 4.2). Next, we use a contrastive selection method to choose the most suitable candidate image for video generation (Section 4.3). This involves encoding both the step description and the visual scene (Section 4.3). Finally, we use the sequential information of the tasks to train our model to accurately select the most coherent image (Section 4.3).
@@ 4.1 Sequential Language Conditioning
Following the work of Bordalo et al. [5], we use an LLM to transform the sequence of text descriptions for each step into visual captions. This has shown to produce Latent Diffusion Model (LDM) prompts that are visually richer, leading to better results [13].
Hence, we train a decoder model $\varphi$ to convert the entire context into one self-contained description,
$\varphi(s_{n}|\{s_{n-1},\ldots s_{1}\}).$ (3)
@@ 4.2 Sequential Denoising Conditioning
Previously, we discussed the sequential dependency assumption ${s_{n}|s_{n-1},\ldots s_{1}}$ for the denoising process input. However, achieving sequential dependency within the denoising process itself poses a challenge. While aligning the input description with the desired output enhances the final result, it does not inherently enforce the generation of visually consistent images. This can lead to accurate depictions of steps but without visual coherence. Therefore, guiding the denoising process is essential to ensure visual coherence across sequential outputs.
We propose a contrastive method to select the image that best represents $s_{n}$ in terms of its description and the preceding scenes. This selective approach strikes a balance between conditioning on all pairs [21], which provides comprehensive information but can be slow and difficult to train, and conditioning on only a single preceding step [5], which offers a faster but less detailed approximation. Our method effectively captures the non-linear nature of the steps in the task that we aim to model. Formally, the denoising iterations follow equation 4, except for the starting iteration $T$ of the reverse diffusion process,
@@
where instead of initializing the latent variable $z_{T}^{i}$ with a random sample from $\mathcal{N}(\mu,\,\sigma^{2})$, we propose to sample denoised latents from prior steps $s_{<n}$ in the sequence. Formally, for each step $n$, we sample $z^{N}_{T}$ from a set of latents produced in previous denoising iterations of earlier steps,
$z_{T}^{n}\sim\{z_{T}^{<n},z_{T-1}^{<n},\ldots,z_{T-w}^{<n}\},$ (6)
where for each step $n$ we consider all previous steps $(n-1)$ and a window of the first $w$ denoising iterations for each step. Finally, each candidate visual scene $v_{n}^{i}$ is then generated with multiple denoising processes using the seeds $\{z_{T}^{<n},z_{T-1}^{<n},\ldots,z_{T-w}^{<n}\}$. By conditioning the generation process for step $n$ on latents from all prior steps in this nonlinear manner, we improve the coherence of the generated sequence.
@@
4.3 Contrastive Selection
We prioritize the visual representation that achieves the best overall consistency across the entire sequence. This selection is achieved by comparing the conditioned scenes on the previous scenes, ensuring the final output maintains visual coherence. To achieve this, we first calculate the dot product between each conditioned scene and the previous scenes, Figure 3. This allows us to measure the similarity between the conditioned scene at step $n$, denoted as $sc_{n}$, and all preceding scenes $sc_{<n}$. We define this similarity as:
$\sum_{k=1}^{n-1}sc_{n}\cdot sc_{k}$ (7)
Next, we apply the softmax function to these similarity scores to convert them into probabilities, making it easier to compare how each conditioned scene is related with the previous scenes. Finally, we select the conditioned scene $sc_{n}^{i}$ with the highest probability to generate the video for the next step. This corresponds to computing the
@@ 4.3 Contrastive Selection
where $\sigma_{\text{CoSeD}}$ is the softmax function applied over all conditioned scenes, and $v_{n}^{i}$ is the image associated with the selected conditioned scene.
By following this process, we ensure that each step in the video is generated based on the conditioned scene that has the strongest visual and contextual relationship to previous steps, optimizing the flow of the overall sequence.
Text and Vision Scene Embeddings
To effectively handle the text and visual modalities of a scene $(s_{n},v_{n})$ in a sequence, we encode both modalities using CLIP [22], see Figure 2. Subsequently, the output of each encoder, denoted as $x\in\mathbb{R}^{t\times d}$, is linearly projected to half $\frac{d}{2}$, where $d$ represents the feature dimension and $t$ is the number of tokens.
For the projection, we use four distinct weight matrices: $W_{IT}$ and $W_{OT}$ for the text embedding, and $W_{IV}$ and $W_{OV}$ for the visual embedding. $W_{IT}$ and $W_{IV}$ project the embeddings of past scenes, while $W_{OT}$ and $W_{OV}$ project the embeddings of the current scene. The projected embeddings are then concatenated into one single vector.
The resulting embedding projections $sc_{n}^{i}$ of all candidate scenes $(s_{n},v_{n}^{i})$ and all past scenes $sc_{<n}$, allow us to represent all scenes within a unified embedding space.
Contrastive Training
Figure 3: Multi-scene V&L contrastive learning uses multiple sequences a cross-entropy loss guides the model to learn the best next scene according to the ground-truth scenes.
During the contrastive selection training phase, we fine-tune a 600K-parameter model to learn the relationships between sequential steps across multiple tasks simultaneously. The model processes a set of $N$ steps (both descriptions and images) from a pool of $M$ tasks. For each task, the model is given a step to be processed in the “next scenes”, while all preceding steps from that task, located in the “past scenes”, serve as context. This setup allows the model to effectively leverage sequential dependencies and learn how each future step relates to its corresponding past steps within the same task, thereby improving coherence (see Figure 3). Formally, we adopted the cross-entropy loss function,
@@
to guide the learning process, by comparing the model’s predictions $\sigma_{CoSeD}(\cdot)$ with one-hot encoded ground truth labels $l_{t,k}$ for each task $t$ and step $k$. These ground truth labels indicate whether a specific step belonged to a task represented in the context. More details can be found in the Annex B.

### v3

@@ Abstract
Generated video scenes for action-centric sequence descriptions, such as recipe instructions and do-it-yourself projects, often include non-linear patterns, where the next video may need to be visually consistent not with the immediately preceding video but with earlier ones. Current multi-scene video synthesis approaches fail to meet these consistency requirements. To address this, we propose a contrastive sequential video diffusion method that selects the most suitable previously generated scene to guide and condition the denoising process of the next scene. The result is a multi-scene video that is grounded in the scene descriptions and coherent w.r.t. the scenes that require visual consistency. Experiments with action-centered data from the real world demonstrate the practicality and improved consistency of our model compared to previous work. Code and examples available at https://github.com/novasearch/CoSeD
@@ 1 Introduction
When people perform tasks involving numerous intricate steps, complementing textual instructions with visual illustrations enhances the user experience [11, 25]. For this reason, various platforms and tools provide multi-scene videos to convey instructional content, such as recipe instructions and do-it-yourself (DIY) projects [16].
State-of-the-art video synthesis methods demonstrate remarkable performance in generating single-scene videos [12, 26, 29, 3, 4]. Yet, only a few works address multi-scene video generation [18, 17, 31]. These methods focus on domains where a single character is central to all scenes, achieving coherence by reusing and combining visual elements across frames. However, multi-scene instructional video synthesis raises a number of challenges. First, the input is a strict sequence of actions, for which it is necessary to generate the full sequence of videos. A model should not generate just the last scene, like GILL [14], and one cannot provide the topic and let the model generate a random sequence of text-video pairs similarly to VideoDrafter [18]. Second, similar to story generation [21, 9, 23, 19], the generative model needs to determine which previous step in the sequence to use as the basis for grounding each new scene. Third, while existing methods are focused on video generation where a single (typically human) character is the center of all scenes [18], instructional videos typically incorporate multiple objects instead of central characters. Hence, we argue that multi-scene instructional video synthesis requires an approach that is sequence-grounded by design, (see Figure 1).
Figure 1: CoSeD is grounded on an input sequence of step actions to synthesize non-linear, multi-scene instructional videos.
To this end, we propose CoSeD (Contrastive Sequential Diffusion learning), a novel approach to instructional video generation. Our method generates candidate images for each step based on the textual description and latent information from previous steps. We then use a contrastive selection approach to choose the best image by evaluating it against prior step descriptions and images. Finally, we use these images to produce a video for each step, ensuring an accurate and consistent representation of the entire task sequence.
CoSeD was able to generate coherent video sequences across diverse instructional content, maintaining high fidelity and relevance in aligning language with vision. CoSeD showed a 20% improvement in human evaluations compared to existing multi-scene methods and was preferred 68% of the time in side-by-side comparisons. Additionally, the compact size of our model enables efficient training and fine-tuning. To summarize our contributions:
@@
  • We introduce contrastive diffusion learning over latents sampled from previous generations.
  • We contribute to a better understanding of the role of seeds and the conditioning of the reverse diffusion process in prior latent representations.
@@ 2 Related Work
Various approaches have been explored to address coherence in image generation. AR-LDM [21] introduces a history-aware autoregressive latent diffusion model that incorporates information from previous steps into the diffusion model’s cross-attention mechanism to guide generation. However, achieving the reported results requires intensive training of the entire pipeline for each dataset, which includes 650 million parameters. Make-a-Story [23] incorporates the complete history of intermediate image representations (latent vectors), which may introduce noise and potentially lead to content generation based on less relevant past information. GILL [14] fuses frozen text-only large language models (LLMs) with pre-trained image encoder and decoder models through a mapping network, enabling multimodal capabilities like image retrieval and generation. However, this can break coherence if retrieved images do not align with the context. SEED-LLaMA [10] integrates a visual tokenizer with a multimodal LLM to process text and images, excelling in multi-turn generation, but struggles with maintaining narrative coherence in story generation tasks.
@@
To generate long single-scene videos Blattmann [4] and Yin [30] generate sparse key frames and interpolate intermediary frames recursively to enhance the frame rate. Extending this idea, Stable Video Diffusion [3] creates a large dataset of annotated video clips by filtering out those with low motion or excessive text, resulting in the generation of higher quality videos. In contrast, Lumiere [2] generates the entire video in a single pass, eliminating the need for sparse key frames and interpolation.
@@
Improvements have also been made in generating coherent multi-scene videos. Video Drafter [18] employs brute-force LLM prompting to create distinct scenes and detailed descriptions for each element. Then it generates image templates that are combined with scene descriptions to produce the final video. Similarly, VideoDirectorGPT  [17] employs a two-stage process in which GPT-4 [20] expands text prompts into detailed descriptions and ensures visual continuity by generating textual descriptions and entity layouts. Mora [31] introduces a multiagent framework, breaking tasks into subtasks like refining prompts [27], and generating images to create small video segments, which are then assembled [7] into a coherent final video, achieving performance comparable to closed-source models such as SORA [6]. However, it relies on using only the last frame of each video segment to start the generation of the next one, which can be problematic if the current step is not directly related to the previous one. StoryDiffusion [32] maintains coherence across frames using a self-attention mechanism and a module for smooth transitions, ensuring that videos faithfully depict the input prompt. Lastly, TALC [1] enhances text-to-video (T2V) models by improving the temporal alignment between video scenes and text segments, improving visual fidelity and narrative coherence.
@@
Building on these advancements, we develop a method that addresses the challenge of maintaining scene coherence while respecting text descriptions, while keeping the model compact to facilitate fine-tuning across multiple domains.
@@ 3 Problem Setting
This section outlines our methodology for generating scene sequences that align with each instruction while preserving continuity with preceding scenes. By enhancing Latent Diffusion Models, we can effectively learn the interdependencies across scenes and guarantee a cohesive visual progression, even when step relations are non-linear.
3.1 Sequential Scene Dependency
@@
In our setting, we depart from the linear dependency assumption used in previous works [31] and acknowledge the possibility of a more complex and non-linear sequential structure [8]. To address this assumption, the model needs to consider not only the current step description and visual scene pair ($s_{n}$, $v_{n}$), but also the pairs from the previous steps and visual scenes, $\{(s_{1},v_{1}),\ldots,(s_{n-1},v_{n-1})\}$. This ensures coherence in the visual elements generated, maintaining consistency, and reflecting the progression of the task, even when individual steps are vague or missing details.
3.2 Sequential Multi-Scene Diffusion
Latent Diffusion Models are designed to synthesize one single image or video at a time. Our goal is to move beyond this limitation and propose a sequential-diffusion model that learns how semantic and visual dependencies should exist in a sequence of multiple scenes. Using the Latent Diffusion Models formulation proposed by [24], the independent denoising process for each isolated step $s_{n}$ of a sequence is the direct application of the model,
@@ 3.2 Sequential Multi-Scene Diffusion
where $z_{t}^{n}$ corresponds to the denoising iteration $t$ of the visual scene $v_{n}$, hence $z_{0}^{n}=E(v_{n})$. Formally, we wish to learn a sequence model that iteratively estimates the $v_{n}$ scene that maximizes the likelihood given the entire sequence of all previous $n-1$ steps. Formally, we have,
@@
where $N$ is the total number of steps in a given sequence. We propose to ensure sequential consistency through the text conditioning encoder $\tau_{\theta}(s_{i})$ and the visual denoising seed $z^{n}_{T}$. The proposed contrastive sequential diffusion learning, Figure 2, incorporates these two methods and will be discussed in the next section.
@@ 4 CoSeD: Contrastive Sequential Diffusion
Our approach aims to find the most accurate image for generating a video that depicts a step of a task. We begin by using a text decoder model to better align the step description with a visual caption/prompt (Section 4.1). Then we generate candidate images based on the information from previous images and the visual caption of the current step (Section 4.2). Next, we use a contrastive selection method to choose the most suitable candidate image for video generation (Section 4.3). This involves encoding both the step description and the visual scene (Section 4.3). Finally, we use the sequential information of the tasks to train our model to accurately select the most coherent image (Section 4.3).
@@ 4.1 Sequential Language Conditioning
Following the work of Bordalo et al. [5], we use an LLM to transform the sequence of text descriptions of each step into visual captions. This has been shown to produce text-to-image prompts that are visually richer, leading to better results [5, 13]. Hence, we train a decoder model $\varphi$ to convert the entire context into one self-contained description,
$\varphi(s_{n}|\{s_{n-1},\ldots s_{1}\}),$ (3)
@@ 4.2 Sequential Denoising Conditioning
Previously, we discussed the sequential dependency assumption $(s_{n}|s_{n-1},\ldots s_{1})$ for the input of the denoising process. However, achieving sequential dependency within the denoising process itself poses a challenge. Although aligning the input description with the desired output enhances the final result, it does not inherently enforce the generation of visually consistent images. This can lead to accurate depictions of steps but without visual coherence. Therefore, guiding the denoising process is essential to ensure visual coherence across sequential outputs.
We propose a contrastive method to select the image that best represents $s_{n}$ in terms of its description and the preceding scenes. This selective approach strikes a balance between conditioning in all pairs [21], which provides comprehensive information but can be slow and difficult to train, and conditioning on only a single preceding step [5], which offers a faster but less detailed approximation. Our method effectively captures the non-linear nature of the steps in the task that we aim to model. Formally, the denoising iterations follow equation 4, except for the starting iteration $T$ of the reverse diffusion process,
@@
where instead of initializing the latent variable $z_{T}^{n}$ with a random sample from $z_{T}\sim\mathcal{N}(\mu,\,\sigma^{2})$, we propose to sample denoised latents from prior steps $s_{<n}$ in the sequence. Formally, for each step $n$, we consider the set of latents produced in previous denoising iterations of earlier steps, i.e.
$\{z_{T}^{i},z_{T-1}^{i},\ldots,z_{T-w}^{i}\}^{i\in(1,\ldots,n-1)},$ (6)
where $i$ indexes all steps from $1$ to $(n-1)$ and $w$ is the window size over the first denoising latents of each step. Finally, all candidate visual scene $v_{n}^{i}$ are generated with the set of latents. By conditioning the generation of step $n$ on latents from all previous steps in this complex non-linear way, the coherence of the generated sequence is improved.
@@
4.3 Multi-Scene Contrastive Selection
Text and Vision Scene Embeddings.
To effectively handle the text and visual modalities of a scene $sc_{n}=(s_{n},v_{n})$ in a sequence, we encode both modalities using CLIP [22], see Figure 2. Subsequently, the output of each encoder, is linearly projected to reduce its dimension to half the original size. For the projection, we use four distinct weight matrices: $W_{IT}$ and $W_{OT}$ for the text embedding, and $W_{IV}$ and $W_{OV}$ for the visual embedding. $W_{IT}$ and $W_{IV}$ project the embeddings of past scenes, while $W_{OT}$ and $W_{OV}$ project the embeddings of the current scene. The projected embeddings are then concatenated into one single vector.
The resulting embedding projections $sc_{n}^{i}$ of all candidate scenes $(s_{n},v_{n}^{i})$ and all past scenes $sc_{<n}$ allow us to represent all scenes within a unified embedding space.
Contrastive Selection.
We prioritize the visual representation that achieves the best overall consistency throughout the sequence. This selection is achieved by comparing the conditioned scenes with the previous scenes, ensuring that the final output maintains visual coherence. To achieve this, we first represent a scene $sc_{n}=(s_{n},v_{n})$ as the concatenation of its text and visual embeddings, and then calculate the dot product between each conditioned scene and the previous scenes, Figure 3. This allows us to measure the similarity between the conditioned scene in step $n$, denoted as $sc_{n}$, and all preceding scenes $sc_{<n}$. We define this similarity as
$\sum_{k=1}^{n-1}sc_{n}\cdot sc_{k}.$ (7)
Next, we apply the softmax function to these similarity scores to convert them into probabilities, making it easier to compare how each conditioned scene is related to the previous scenes. Finally, we select the conditioned scene $sc_{n}^{i}$ with the highest probability of generating the video for the next step. This corresponds to computing the
@@ 4.3 Multi-Scene Contrastive Selection
where $\sigma_{\text{CoSeD}}$ is the softmax function applied over all conditioned scenes, and $v_{n}^{i}$ is the image associated with the selected conditioned scene $sc^{i}_{n}=(s_{n},v^{i}_{n})$.
By following this process, we ensure that each step in the video is generated based on the conditioned scene that has the strongest visual and contextual relationship to the previous steps, optimizing the flow of the overall sequence.
Contrastive Training.
Figure 3: Multi-scene V&L contrastive learning uses multiple sequences. This multi sequence information serves as both positive and negative pairs helping the model to learn the best next scene according to the ground-truth scenes.
During the contrastive selection training phase, we fine-tune a 600K-parameter model to learn the relationships between sequential steps across multiple tasks simultaneously. The model processes a set of $N$ steps (both descriptions and images) from a pool of $M$ tasks. For each task, the model is given a step to be processed in the ’next scenes’, while all preceding steps of that task, located in the ’past scenes’, serve as context. This setup allows the model to effectively leverage sequential dependencies and learn how each future step relates to its corresponding past steps within the same task, thus improving coherence (see Figure 3). Formally, we adopted the cross-entropy loss function,
@@
to guide the learning process, by comparing the model’s predictions $\sigma_{CoSeD}(\cdot)$ with one-hot encoded ground truth labels $l_{t,k}$ for each task $t$ and step $k$. These ground truth labels indicate whether a specific step belonged to a task represented in the context. More details can be found in the appendix file.